ISS Daily Summary Report – 07/25/16

Biological Research In Canisters (BRIC) 23: Over the weekend, the crew transferred all four of the BRIC-23 canisters that were actuated last week to the Minus Eighty Degree Celsius Laboratory Freezer for ISS (MELFI).  Two of the four canisters were transferred to MELFI following a 24-hour growth period and the remaining two canisters were transferred following a 48-hour growth period. The BRIC-23 investigation studies Bacillus subtilis spores and Staphylococcus aureus cells to understand how they respond to the stressful environment of space. Results from this investigation improve the understanding of how microbes adapt to spaceflight, including whether their adaptations change antibiotic effectiveness, which benefits efforts to maintain crew member health.

NanoRacks Module 9: Over the weekend the crew completed the second session of the NanoRacks Module-9 experiment by activating, shaking, and deactivating the mixture tubes. The NanoRacks Module-9 experiment is a collection of student research projects utilizing the NanoRacks Mixsticks. Student teams from across the United States design their own experiments using flight approved fluids and materials. The investigation consists of several science experiments flown in a NanoRacks Module on board the ISS.

Biological Rhythms 48 Holter Start: Earlier this morning the crew continued the operations phase of the Biological Rhythms experiment by attaching the Digital Walk Holter Electrocardiogram (ECG) and electrodes. The objective of the Japan Aerospace Exploration Agency (JAXA) Biological Rhythms 48 is to study the effects of long-term microgravity exposure on heart function by analyzing an astronaut’s electrocardiogram for 48 hours.

Mouse Epigenetics Maintenance Operations: The crew refilled water in the Mouse Habitat Cage Unit located in the Cell Biology Experiment Facility (CBEF). The Mouse Epigenetics investigation studies altered gene expression patterns in the organs of male mice that spend one month in space, and also examines changes in the deoxyribonucleic acid (DNA) of their offspring. Results from the investigation identify genetic alterations that happen after exposure to the microgravity environment of space. 

Fluid Shifts: The crew conducted the first Fluid Shifts operation of this session by configuring the Refrigerated Centrifuge for sample load operations, performing body sample collections and stowing the samples in a MELFI. The crew also collected a galley water sample and stowed in the MELFI prior to ingestion of a Tracer solution from the Tracer Syringe. Fluid Shifts is a joint USOS Russian experiment that measures how much fluid shifts from the lower body to the upper body, in or out of cells and blood vessels, and determines the impact these shifts have on fluid pressure in the head, changes in vision and eye structures.

Meteor Hard Drive Change-out: The crew removed and replaced the hard drive in the Meteor Laptop located in the Window Observational Research Facility (WORF). The investigation provides the first space-based measurement of meteor flux as well as monitoring of carbon-based compounds. Continuous measurement of meteor interactions with the Earth’s atmosphere could also spot previously unforeseen meteor showers. 

Habitability Human Factors Directed Observations: The crew recorded and submitted a walk-through video documenting observations of life onboard ISS, providing insight related to human factors and habitability. The Habitability investigation collects observations about the relationship between crew members and their environment on the ISS. Observations can help spacecraft designers understand how much habitable volume is required, and whether a mission’s duration impacts how much space crew members need.

“Road To” Extravehicular Activity (EVA): This week begins preparation tasks for the IDA2 EVA currently scheduled for August 18. Today the crew was scheduled to complete the following EVA Mobility Unit (EMU) tasks, but only completed the first three due to scheduling constraints:

  • Remove EMU 3008 from the Fwd EMU Don/Doff Assembly (EDDA) and stow in the Crew Lock (CL).
  • Remove EMU 3006, which was delivered on SpX-9, from the SEMU Launch Enclosure (SLE), install on the Fwd EDDA and configure for on-orbit operations.
  • Configure and install EMU 3005 in the SLE for return on SpX-9.
  • Retrieve EMU 3003 from the CL and install on Aft EDDA. [Deferred]
  • Remove EMU 3006 from Fwd EDDA and install EMU 3008. [Deferred]

Potable Water Dispenser Filter Maintenance: The crew R&Rd the PWD filter and cleaned the fan filter areas behind the PWD. This preventive maintenance is required every 18 months.

Transition to the Power Management Control Application (PMCA) and Photovoltaic Control Application (PVCA) R5 (PMPV R5) software suite – The X2 PMPV R5 software transition began over the weekend and will continue into Tuesday.  PMCA and PVCA R5 contains new software to manage the Li-Ion batteries that will arrive on HTV-6. In addition to the PMCA and PVCA updates, there also are updates to the Command and Control (C&C) and Portable Computer System (PCS) software.  All software has good health signatures to date.

Today’s Planned Activities
All activities were completed unless otherwise noted.

  • FS – saliva sampling
  • FS – urine sampling
  • FS – install cooling centrifuge
  • FS – blood sampling
  • FS – water sampling from on-board kitchen and acceptance of radiopaque marker
  • FS – configure cooling centrifuge
  • CASCADE. Manual mixing in bio reactor / r/g 2888
  • MOUSE- prepare equipment for experiment
  • FLUID SHIFTS. Close out ops with cooling centrifuge
  • FLUID SHIFTS. Placement of samples into MELFI
  • MOUSE- clean the module
  • FS- urine sampling
  • PILOT-Т. Preparation for experiment r/g 2885
  • WRS – water sampling
  • Crew handover / РПС
  • BLR48 – start of measurements
  • FLUID SHIFTS. Placement of samples into MELFI
  • Recording of audio gram with EARQ s/w
  • PILOT-Т. Experiment run r/g 2885
  • XF305 video camera adjustment
  • MOUSE- replacement of water in the module
  • EMU SS backpack replacement
  • SLE – rotation, part 1
  • EMU-reconfiguration after retrieval from transportation container
  • FS – activation of cooling centrifuge
  • FS – blood sampling
  • FS – sampling
  • FS – configure cooling centrifuge
  • FLUID SHIFTS. Установка проб в MELFI
  • Cargo ops in Progress №433 (on DC1) and IMS update r/g 2834, 2835
  • PILOT-Т. Experiment run. Tagup with specialists. / r/g 2886
  • Activity for deployment of FMK monitors
  • HMS – eye view check
  • Sampling in GSC container
  • FLUID SHIFTS. Close out ops with cooling centrifuge
  • Testing of vision functions – questionnaire fill out
  • EMU SS – SEMU short EMU – container for delivery to orbit (SLE) – preparation for return
  • FLUID SHIFTS. Placement of samples into MELFI
  • Transfusion through micro biological filter
  • Cargo ops in Progress №433 (on DC1) and IMS update r/g 2834, 2835
  • PILOT-Т. Close out ops. Tagup with specialists. / r/g 2886
  • FS – sampling
  • FLUID SHIFTS. Placement of samples into MELFI
  • SOZh maintenance
  • FS – urine sampling
  • HMS – eye view check
  • Preparation for replacement of unit 800А (А74) on FGB
  • Ham radio pass from Columbus
  • FLUID SHIFTS. Placement of samples into MELFI
  • Testing of vision functions – questionnaire fill out
  • ARED – clean flywheel cylinder
  • HABIT- preparation for experiment
  • 24-hour long registration of arterial pressure (start) / r/g 2892
  • SM ventilation equipment preventive maintenance/ Group A
  • SLE – rotation, part 2
  • MRF filter air washing
  • DRAGON. Cargo transfer
  • PWD dispenser – filter R&R
  • 24-hour  long ECG registration (start) / r/g 2893
  • Update anti-virus base on [ВКС] laptops
  • DRAGON. Cargo transfer
  • EMU swap- [Deferred]
  • Cargo ops in Progress №433 (on DC1) and IMS update r/g 2834, 2835
  • TOCA – data entry
  • PR2 – install equipment
  • WRS. Drain waste tank (start)
  • Dragon – cargo transfer conference
  • IMS update
  • CONTROL. Change measurement mode of INDICATOR ISS science equipment r/g 2884
  • Replace hard drive in METEOR laptop
  • WRS. Drain waste tank (term)
  • ВЕЛО-1 exercise
  • BIOPLENKA. Install thermostat CRYOGEM-03  / r/g 2891
  • BIOPLENKA. Activate thermostat CRYOGEM-03  in + 37 deg. С mode/ r/g 2891
  • MRF filter air washing
  • D/l physical exercise data via  ОСА
  • CASCADE. Manual mixing in bio reactor / r/g 2888
  • Water transfer into tank through MRF filter (term)
  • BIOPLENKA. Secure biofilm of CONSTANTA №3-1 cassette and place in ТБУ-В №04 r/g 2890
  • BIOPLENKA. Imagery during placement into thermostat / r/g 2889
  • Preparation for anti-virus scanning of [ВКС] laptops 

Completed Task List Items

  • WHC KTO seat R&R
  • JAXA video take Pt 2 [Active]
  • Dragon transfer ops [Active]

Ground Activities
All activities were completed unless otherwise noted.

  • Cargo transfer support
  • EVA tasks support
  • Nominal ground commanding.

Three-Day Look Ahead:

  • Tuesday, 07/26: SAM removal from JEMAL ST, PEPS inspection, Fluid Shifts
  • Wednesday, 07/27: Fluid Shifts, Mouse cage maintenance, Fine Motor Skills, Airway Monitoring hardware locate
  • Thursday, 07/28: Fluid Shifts, Mouse cage maintenance, Airway Monitoring hardware deploy, EMU resize, Active Dosimeter install/checkout

QUICK ISS Status – Environmental Control Group:







[СКВ] 1 – SM Air Conditioner System (“SKV1”)


[СКВ] 2 – SM Air Conditioner System (“SKV2”)


Carbon Dioxide Removal Assembly (CDRA) Lab


Carbon Dioxide Removal Assembly (CDRA) Node 3


Major Constituent Analyzer (MCA) Lab


Major Constituent Analyzer (MCA) Node 3


Oxygen Generation Assembly (OGA)


Urine Processing Assembly (UPA)


Trace Contaminant Control System (TCCS) Lab


Trace Contaminant Control System (TCCS) Node 3

Full Up